News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Sunday 5 September 1999 9.00pm EST

During the last 48 hours a series of HRC-S focus measurements were completed successfuly. Measurements included a series of shutter scans that provided focus centroid data. The CXC science team completed analysis of the data and provided the new SIM position for HRC-S in focus in preparation for the deployment of the Low Energy Transmission Grating (LETG).

The LETG deployment procedure was conducted between 6PM and 9PM on 9/5 and the team successfully released both launch locks and inserted the LETG with HRC-S in the focus position. The science team confirmed that data from Capella looks good and that initial assessment indicates that the LETG is functioning as expected. Great news that the final mechanism has been deployed on Chandra! The LETG will be left inserted for 11 hours and then retracted via hardware command.

A telecon was held to discuss the observed B-side voltage during HETG insertion and retractions. The decision was made to continue with the LETG deployment and set a relay to allow monitoring of B-side telemetry during insertion and retraction. Expectations were that simlar voltage (probably caused by back-EMF) would be seen with LETG. This was confirmed during the procedure with a value of 6.9 volts compared with 6.0 volts seen with HETG. Further analysis will continue to confirm that the voltage is due to back-EMF.





Return to Status Reports